OpenStudy (elise_a18):

Check please--- Geometric Sequences hat is the sum of the first five terms of the geometric sequence in which a 1=20 and r=1/4

OpenStudy (elise_a18):

I got 164.1406

OpenStudy (afro):

aren't you suppose to multiply the first by 1/4 which is the rate?

OpenStudy (elise_a18):

I don't think that's the first step, I think raising 1/4 to the power of 5 is the first

OpenStudy (afro):

O_O why would you do that when you could just multiply 1/4 by 20

OpenStudy (afro):

and do it with each answer you get until you get to 5

OpenStudy (elise_a18):

because I'm finding the partial sum of the first 5 terms

OpenStudy (afro):

is that what the question asked? or are you suppose to find the fifth term in the rule of geometric sequences

OpenStudy (elise_a18):

I need to find the Sum of all the terms up until the fifth

OpenStudy (afro):

yea, well that's quicker, but how did you get that huge number

OpenStudy (afro):

multiply that decimal number to the 20

OpenStudy (elise_a18):

instead of using just .25, i used 1.25 and raised it to the 5th power before subracting it from 1

OpenStudy (elise_a18):

i was trying to learn from my last problem... :/

OpenStudy (afro):

something totally different the last problem you were dealing with a different formula

OpenStudy (elise_a18):

I just don't know which formula to use... i did it again with another and got 26.64025

OpenStudy (elise_a18):

@agent0smith @zarkam21

OpenStudy (elise_a18):

@zepdrix does 26.64.... seem right to you?

zepdrix (zepdrix):

Probably >.< I dunno I'm trying to remember the formula thingy

OpenStudy (elise_a18):



zepdrix (zepdrix):

I worked it out on paper, I think it's this for partial sum of geometric,\[\large\rm S_n=a_1\cdot\frac{1-r^n}{1-r}\]

zepdrix (zepdrix):

\[\large\rm S_5=20\cdot\frac{1-r^5}{1-r}\]

zepdrix (zepdrix):

Plugging in our rate as well,\[\large\rm S_5=20\cdot\frac{1-.25^5}{1-.25}\]

zepdrix (zepdrix):

Which works out to 26.640625 yay good job \c:/

OpenStudy (elise_a18):

yee, ok then it says to express the answer as an improper fraction

zepdrix (zepdrix):

Oh ok. Then I guess we'll have to leave the calculator out of this. Back up to this point,\[\large\rm S_5=20\cdot\frac{1-.25^5}{1-.25}\]Use 1/4 instead of the decimal value,\[\large\rm S_5=20\cdot\frac{1-\frac14^5}{1-\frac14}\]And simplify from there.

OpenStudy (elise_a18):

i don't know how u_u

OpenStudy (agent0smith):

I'm sure you can simplify some of it... at least the (1/4)^5 and the 1 - 1/4